A car slows from 22 m/s to 3.0 m/s with a constant acceleration of -2.1 m/s2.  How long does it require in the unit of seconds?


1
Expert's answer
2021-10-07T08:49:37-0400

By definition, the acceleration is given as follows:


"a = \dfrac{v_f-v_i}{t}"

where "v_f = 3m\/s,v_i = 22m\/s" are final and initial speeds respectively, and "t" is the time. Expressing the time and substituting "a = -2.1m\/s^2", obtain:


"t = \dfrac{v_f-v_i}{a} = \dfrac{3m/s - 22m/s}{-2.1m/s^2} \approx 9.0s"

Answer. 9.0 s.